OpenStudy (anonymous):

After solving this, exponential equation I get two answers for x, my question is are both he values valid? Or is any one value valid?

OpenStudy (anonymous):

depends on the equation and your solution.

OpenStudy (anonymous):

\[4^{x ^{2}-6} - 16 ^{x+1}\]

OpenStudy (anonymous):

This is not an equation.

OpenStudy (anonymous):

sorry by mistakenly put - instead of = @FoolForMath

OpenStudy (anonymous):

@ash2326 Need help!

OpenStudy (anonymous):

^^^ The whole expression equal to zero.

OpenStudy (saifoo.khan):

call 15 for help.

OpenStudy (saifoo.khan):

\[\Large 4^{x^2 - 6} - 4^{2(x+1)} = 0\]

OpenStudy (anonymous):

^ Yes I know my question is After solving this, exponential equation I get two answers for x, my question is are both he values valid? Or is any one value valid?

OpenStudy (anonymous):

Okay then \[ x^2 - 6 = 2x+2 \implies x= -2\; or \; x=-4 \]

OpenStudy (saifoo.khan):

Try inserting them back in the place of "x" one by one.

OpenStudy (anonymous):

@saifoo.khan But that can be time consuming, (specially in SAT, where you're bound in terms of time) ...

OpenStudy (anonymous):

I just want to ask Whenever you get such type of exponential equation to solve, and you get two answers for x, are both valid? Or how to determine ?

OpenStudy (saifoo.khan):

No. like let's say you get 2. insert 2 back in there and if you get that equal to 0.. then it means the solution is correct.

OpenStudy (anonymous):

Or say you have something with \( e^x \) here x>0 (always). I can't think of any thing bad happening other wise.

OpenStudy (anonymous):

Rephrasing my earlier comment, if for instance you get an imaginary root while solving the quadratic then that is to be eliminated.

OpenStudy (ash2326):

@Zeerak you understood?